OpenStudy (anonymous):

Solve the absolute value inequality: |5x + 3| ≥ 13 x ≤ -3.2 and x ≥ 2 x ≥ -3.2 or x ≤ 2 x ≤ -3.2 or x ≥ 2 x -3.2 or x ≥ 3.2

OpenStudy (anonymous):

\[|x|>a\], with a>0, is the same as x>a or x<-a so rewrite your question like this: \[5x+3 \ge 13 or 5x+3 \le -13\] and solve for x in both parts of the disjunction

OpenStudy (anonymous):

@d im looking for a complete example so i could do the rest on my own thnks

OpenStudy (anonymous):

ok 5x+3>=13 5x>=10 x>=2 5x+3<=-13 5x<=-16 x<=-16/5 So, x <=-16/5 or x>=2

OpenStudy (anonymous):

-16/5=-3.2

OpenStudy (anonymous):

so the first one just to clarify? @d

OpenStudy (anonymous):

no, both. the solution set is {x|x<=-3.2 or x>=2}

OpenStudy (anonymous):

ah the third one!

OpenStudy (anonymous):

to remove the absolute value signs, you write a disjunction like I did